An investment offers to pay $500 a year forever starting at the end of year 5. If the interest rate is 10%, what is the investment's value today? (Assume annual compounding)

  • $3756.57

  • $3722.43

  • $3415.07

  • $3104.61
Answer #1

Present value at year 4 = Perpetual cash flow / interest rate

Present value at year 4 = 500 / 0.1

Present value at year 4 = 5,000

Value today = Present value / (1 + r)n

Value today = 5,000 / (1 + 0.1)4

Value today = 5,000 / 1.4641

Value today = $3,415.07